11.A college professor wants to report and estimate the mean weekly salaries of his part-time working students. A sample of 7 students who work part-time showed the following weekly wages: $107 $192 $97 $295 $105 $151 $91 Construct a confidence interval for the population mean with a 90% confidence level? A. [79.77, 216.80] B. (93.87, 202.70 C. [143.93, 152.64] D. 199.43, 105.68]

12. You want to open a NACHO BIZNESS Mobile Food TruckNOVA. A study would be needed to estimate the proportion of students who would choose to buy nachos for lunch over EuroCafe. You want to be within 5% the proportion of students who would opt for the food truck. A 95% degree of confidence is to be used. You estimate that the proportion of students who would eat nachos for lunch is 20%. HowmanvstudentsNOVAshouldbasurveyed? A. 288 B. 289 C. 389 D. 390
Answer #1

Ans:

1)n=7

df=7-1=6

critical t value=tinv(0.1,6)=1.943

sample mean=148.29

sample standard deviation,s=74.087

Margin of error=1.943*74.087/sqrt(7)=54.41

lower limit=148.29-54.41=93.88

upper limit=148.29+54.41=202.70

Option B is correct.(93.87, 202.70)